# Impact of 3D macro physics and nuclear physics on the $p$ nuclei in $\mathrm{O}$-$\mathrm{C}$ shell mergers

### v1

@@ 3.3 Impact from varying the ingestion rate
Here we present the impact of entraining $\mathrm{C}$-shell material with $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ as well as no entrainment as explained in Section 2.2. Figure 11 shows the results for the MLT mixing scenario and Figures B1B4 in Appendix B show the results for all downturn scenarios.
@@
Figure 11: The overproduction compared to initial of the $p$ nuclei for the MLT mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.22\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount.
@@
The results show that the production of the $p$ nuclei is monotonically increased by the ingestion of $\mathrm{C}$-shell material for all isotopes except $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$ who exhibit the opposite behaviour for all mixing scenarios. The only exception is that $\mathrm{{}^{74}Se}$ production increases for the 10×10$\times$10 × and 50×50$\times$50 ×3D-inspired scenarios with ingestion rate. This is because with higher ingestion rates, the more stable isotopes enter the shell as demonstrated in Section 3.1. The lightest three have the opposite behaviour because without ingestion they are destroyed less by $(n,\gamma)$ reactions. The difference in $\mathrm{OP}$ between two ingestion rates is largely uniformly for $\mathrm{{}^{92}Mo}$-$\mathrm{{}^{196}Hg}$. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}$ between the different ingestion rates for the MLT and 3D-inspired scenarios is $1.22$, $1.58$, $1.64$, $1.78$, and $1.84\,\mathrm{dex}$. This shows that the non-linear and non-monotonic behaviour in Section 3.2 is because of the changing location of peak burning in the convective-reactive environment and not just more material being present.
@@
Without ingestion of $\mathrm{C}$-shell material, the MLT, 1×1$\times$1 ×, and 3×3$\times$3 × scenarios see little to no production, but for the 10×10$\times$10 × and 50×50$\times$50 × scenarios there is a significant underproduction of the heavier $p$ nuclei. This is because in the fastest cases the species are advected to the bottom of the shell where the temperatures are purely destructive and there is no replenishment from the $\mathrm{C}$-shell. This shows that entrainment of $\mathrm{C}$-shell material is necessary for significant contributions from pre-explosive $\gamma$ process.
@@
The sign of the ratio of isotopic pairs is largely preserved across the ingestion rates for the MLT, 1×1$\times$1 ×, 3×3$\times$3 ×, and 50×50$\times$50 × scenarios, but the magnitude of the ratio does change. The few exceptions are $\mathrm{{}^{162,164}Er}$ in the 1×1$\times$1 × and 3×3$\times$3 × scenarios who become roughly equal for the fastest ingestion rate, $\mathrm{{}^{115}Sn}$ which becomes more abundant than $\mathrm{{}^{112,114}Sn}$ in the 3×3$\times$3 × scenario. In the 10×10$\times$10 × scenario production of isotopic pairs is largely equal for the slower ingestion rates, but at the fastest ingestion rate the ratio can become unequal. This shows that the the ingestion rate matters for the magnitude of the ratio between isotopic pairs, but that the sign of the ratio depends on which diffusion profile is used.
3.4 Impact from dips from convective quenching
@@
Here we present the impact of dips from convective quenching using the profiles shown in Figure 5 from Section 2.2 which represent GOSH-like feedback and a partial merger. The results are shown in Figure 12.
@@
Figure 12: The overproduction compared to initial of the $p$ nuclei for the MLT, GOSH-like, and partial merger scenarios with ingestion rate $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.51\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
@@
The dip functions as a barrier for the convective-reactive flow, and can section off parts of the $\mathrm{O}$-shell. The partial merger scenario both limits the ingested $\mathrm{C}$-shell material and slows down the material from deeper in the $\mathrm{O}$-shell from advecting up to the top of the shell where very few reactions occur. The GOSH-like dip slows down the material from reaching their preferential temperatures, and also prevents material at the deepest part of the $\mathrm{O}$-shell from mixing up to the top of the shell which keeps it at hotter temperatures where it can be destroyed. This is also why the deeper dips suppress production more, as the velocities are slower by an additional factor of $10$ at the deepest point of the dip.
The suppression of production is largely uniform for all isotopes except for $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, $\mathrm{{}^{84}Sr}$, and $\mathrm{{}^{180}Ta}$. Additionally, it appears that isotopes $A\geq 138$ are more strongly affected by the stronger GOSH-like dip. The isotopes $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$ have a minor increase from these dips because the location of their production is at the bottom of the $\mathrm{O}$-shell where the dips are not present. The boost in production of $\mathrm{{}^{180}Ta}$ is because the peak production and destruction locations happen to be centered at the exact same location as the GOSH dip, which lowers its destruction and slightly boosts its production. Isotopic ratios are not significantly affected by the dips, although the magnitude of the ratios can increase slightly. This demonstrates the importance of both the location and magnitude of the convective dips in the $\mathrm{O}$-shell for convective-reactive $\gamma$ process.
3.5 Nuclear physics impact and mixing dependencies
@@
The MLT mixing scenario has an average spread in production of $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.56\;\mathrm{\,\mathrm{dex}}$, and the 3D-inspired scenarios have an average spread of $0.59\;\mathrm{\,\mathrm{dex}}$, $0.69\;\mathrm{\,\mathrm{dex}}$, $0.76\;\mathrm{\,\mathrm{dex}}$, and $0.79\;\mathrm{\,\mathrm{dex}}$ for the 1×1$\times$1 ×, 3×3$\times$3 ×, 10×10$\times$10 ×, and 50×50$\times$50 × scenarios respectively. We find that the spread in production increases with mixing speed because the material is able to reach hotter temperatures and the number of possible nucleosynthetic pathways changes.
@@
The spread of an individual isotope is dependent on the mixing scenario. Species like $\mathrm{{}^{106}Cd}$, $\mathrm{{}^{156}Dy}$, and $\mathrm{{}^{180}W}$ have a different spread in production for each mixing scenario. While the change can be monotonic for species like $\mathrm{{}^{180}W}$, $\mathrm{{}^{184}Os}$, and $\mathrm{{}^{190}Pt}$, for $\mathrm{{}^{106}Cd}$ and $\mathrm{{}^{112}Sn}$, and $\mathrm{{}^{130}Ba}$ they decrease for the 50×50$\times$50 ×3D-inspired scenario.
@@
It is clearly important to consider the mixing conditions if an experiment is to be proposed to measure a reaction rate. As Table 2 shows for the 50×50$\times$50 ×3D-inspired scenario, there are many reactions even for a single isotope that can be correlated uniquely in that mixing scenario. This clearly shows that a decreasing radial velocity profile and the exact magnitude of the mixing speeds are crucial for understanding the nuclear reactions in this convective-reactive environment.
There are also correlated reactions that all mixing scenarios share as shown in Table 3. Additionally, all downturn cases share correlations not found in the MLT scenario: $X(\mathrm{{}^{115}Sn})$ with $\mathrm{{}^{110}Sn}(\gamma,\alpha)$ and $X(\mathrm{{}^{138}Ce})$ with $\mathrm{{}^{138}Nd}(\gamma,p)$. However, the shared correlations are not of equal strength across all mixing scenarios. As an example the final mass fraction of $\mathrm{{}^{156}Dy}$ is correlated with $\mathrm{{}^{160}Er}$$(\gamma,\alpha)$, but in the 50×50$\times$50 ×3D-inspired the correlation is much weaker. This underscores the possible difficulties in using 1D astrophysical sites to identify important reactions for nuclear physics experiments.
@@
${}^{144}\mathrm{Sm}$ ${}^{196}\mathrm{Pb}(\gamma,n)$  ${}^{142}\mathrm{Sm}(\gamma,n)$  ${}^{142}\mathrm{Sm}(\gamma,p)$  ${}^{143}\mathrm{Sm}(\gamma,n)$
@@
The spread in production for varying our adopted nuclear physics rates for each of the mixing scenarios is comparable to the spread seen from varying the mixing conditions. The results in Section 3.4 have a spread of $0.51\;\mathrm{\,\mathrm{dex}}$, Section 3.2 $0.96\;\mathrm{\,\mathrm{dex}}$, and the maximum spread in Section 3.3 is $1.84\;\mathrm{\,\mathrm{dex}}$. The average spread in production from varying the nuclear physics rates ranges from $0.56$-$0.79\;\mathrm{\,\mathrm{dex}}$.
Rauscher et al. (2016) studied the impact of nuclear uncertainties for the explosive $\gamma$-process for a $15\;\mathrm{\mathrm{M}_{\odot}}$ and two $25\mathrm{M}_{\odot}$ models with solar metallicity using a Monte Carlo method. The spread in their 90% probability interval for the $15\;\mathrm{\mathrm{M}_{\odot}}$ model was $0.61\;\mathrm{\,\mathrm{dex}}$, $0.63\;\mathrm{\,\mathrm{dex}}$ for the $25\mathrm{M}_{\odot}$ KEPLER model, and $0.99\;\mathrm{\,\mathrm{dex}}$ for the model from Hashimoto et al. (1989). This is comparable to the maximum spread found in this work, although we are considering pre-explosive $\gamma$ process. Rauscher et al. (2016) also provide tables of correlated rates, but only a handful of these rates appear in our Tables C1C5 including those they find with $r_{\mathrm{P}}\geq|0.65|$. This is because the convective-reactive environment we consider allows for different reaction pathways to be favoured depending on the mixing conditions.
4 Discussions and Conclusion
In this paper, we have shown that understanding the mixing details in the $\mathrm{O}$-burning shell during an $\mathrm{O}$-$\mathrm{C}$ shell merger is crucial for accurately modelling the nucleosynthesis of the $p$ nuclei. This work raises the importance of 3D hydrodynamic simulations for understanding the nucleosynthesis in O-C shell mergers (Bazan & Arnett, 1994; Yadav et al., 2020; Rizzuti et al., 2024). We have demonstrated the convective-reactive nature of the nucleosynthesis in the $\mathrm{O}$-shell, where the timescales for advection and reaction are comparable and our results show that:
@@ 4 Discussions and Conclusion
  • The ratio of isotopic pairs are sensitive to the mixing scenario.
@@
Although we find that the mixing conditions significantly impact the results of Section 3.5, not all scenarios are equally likely to represent the conditions in a merger. The MLT mixing scenario, 1×1$\times$1 ×, and 50×50$\times$50 × case may not be representative of the conditions for realistic O-C shell mergers. 3D hydrodynamic simulations show the $\mathrm{O}$-shell has a downturn to radial convective velocities and mixing speeds roughly $3-10$ times larger than what MLT predicts (Jones et al., 2017; Andrassy et al., 2020; Rizzuti et al., 2024). This suggests that the 3×3$\times$3 × and 10×10$\times$10 × downturn mixing scenarios are likely more representative of the conditions in a merger. The exploration done in this paper shows the importance of understanding the mixing conditions, but not all scenarios are equally likely to represent the conditions in a merger.
The results in this work are important for the interpretation of presolar grains. Fok et al. (2024) argue that the nucleosynthesis in $\mathrm{O}$-$\mathrm{C}$ shell mergers matter when interpreting the ${}^{29}\mathrm{Si}/^{28}\mathrm{Si}$ ratio seen in grains. As shown in this work, the ratio of isotopic pairs is sensitive to the mixing conditions in the $\mathrm{O}$-shell. This means that comparing the grains data to the results of this work can be used as a diagnostic tool to constrain the mixing details of $\mathrm{O}$-$\mathrm{C}$ shell mergers and connect measured isotopic ratios to 3D hydrodynamics.
There are limitations to the results provided in this work and further extensions that could be done. We have focused on the impact of mixing conditions and varying reaction rates, but do not investigate how different stellar conditions are relevant to the nucleosynthesis or if changing the stable seeds from the $\mathrm{C}$-shell impact these results. The distribution of stable seeds and the metallicity of the star are of significant importance for the nucleosynthesis of the $p$ nuclei (Travaglio et al., 2015; Battino et al., 2020). It is possible that earlier in stellar evolution that weak $s$ process in the $\mathrm{C}$-shell could modify the stable seeds (Pignatari et al., 2010), which could be relevant for the lighter $p$ nuclei.
@@
Finally, $\mathrm{O}$-$\mathrm{C}$ shell mergers are crucial to the nucleosynthesis of a massive star prior to the CCSN regardless of explosive energy (Roberti et al., 2024). Even if the results in Figure 14 do not represent the whole of the nucleosynthesis of the $p$ nuclei, they are still crucial for understanding the nucleosynthesis in the $\mathrm{O}$-shell prior to the CCSN.
@@
The results of this work have implications beyond the $p$ nuclei. The light odd-Z elements $\mathrm{P}$, $\mathrm{Cl}$, $\mathrm{K}$, and $\mathrm{Sc}$ are also produced during O-C shell mergers and based on our preliminary results are likewise impacted by varying the mixing conditions (Ritter et al., 2018a; Roberti et al., 2025). The long-lived radioactive isotope $\mathrm{{}^{40}K}$ ($t_{1/2}=1.25\times 10^{9}\;\mathrm{\mathrm{yr}}$), which is relevant to the heating of planets early in their formation (Frank et al., 2014; O’Neill et al., 2020), is also affected along with the stable $\mathrm{K}$ isotopes $\mathrm{{}^{39}K}$ and $\mathrm{{}^{41}K}$. Finally, observations have found $\mathrm{P}$-enhanced stars (Masseron et al., 2020; Brauner et al., 2023, 2024) which could be explained by a $\mathrm{O}$-$\mathrm{C}$ shell merger from a previous massive star.
@@
O-C shell mergers have a potentially huge impact on galactic chemical evolution models (Ritter et al., 2018a), and massive star models show this feature regardless of metallicity and stellar evolution model (Roberti et al., 2025). Further work is needed to understand the impact of the macrophysical uncertainties in 1D stellar models on the nucleosynthesis of these light odd-Z elements.
@@
The authors would like to thank Stephen Mojzsis for the useful discussions. MP acknowledges the support to NuGrid from the “Lendulet-2023” Program of the Hungarian Academy of Sciences (LP2023-10, Hungary), the ERC Synergy Grant Programme (Geoastronomy, grant agreement number 101166936, Germany), the ERC Consolidator Grant funding scheme (Project RADIOSTAR, G.A. n. 724560, Hungary), the ChETEC COST Action (CA16117), supported by the European Cooperation in Science and Technology, and the IReNA network supported by NSF AccelNet (Grant No. OISE-1927130). MP also thanks the support from NKFI via K-project 138031 (Hungary). FH is supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant and acknowledges support from the NSERC award SAPPJ-797 2021-00032 Nuclear physics of the dynamic origin of the elements. We acknowledges support from the ChETEC-INFRA project funded by the European Union’s Horizon 2020 Research and Innovation programme (Grant Agreement No 101008324). This research has used the Astrohub online virtual research environment (https://astrohub.uvic.ca), developed and operated by the Computational Stellar Astrophysics group (http://csa.phys.uvic.ca) at the University of Victoria and hosted on the Computed Canada Arbutus Cloud at the University of Victoria. This work benefited from interactions and workshops co-organized by The Center for Nuclear astrophysics Across Messengers (CeNAM) which is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-SC0023128.

### v2

@@ III.3 Impact from varying the ingestion rate
Here we present the impact of entraining $\mathrm{C}$-shell material with $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ as well as no entrainment as explained in Section II.2. Figure 11 shows the results for the MLT mixing scenario and Figures B1B4 in Appendix B show the results for all downturn scenarios.
@@
Figure 11: The overproduction compared to initial of the $p$ nuclei for the MLT mixing scenario for no ingestion, $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=1.22\;\mathrm{\,\mathrm{dex}}$ excluding the no ingestion case. $\mathrm{OP}=0$ is the initial amount. Average $\mathrm{OP}$ for each scenario are provided as dashed lines and corresponds to the values presented in Table 1.
@@
The results show that the production of the $p$ nuclei is monotonically increased by the ingestion of $\mathrm{C}$-shell material for all isotopes except $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$ who exhibit the opposite behaviour for all mixing scenarios. The only exception is that $\mathrm{{}^{74}Se}$ production increases for the $10\times$ and $50\times$3D-inspired scenarios with ingestion rate. This is because with higher ingestion rates, the more stable isotopes enter the shell as demonstrated in Section III.1. The lightest three have the opposite behaviour because without ingestion they are destroyed less by $(n,\gamma)$ reactions. The difference in $\mathrm{OP}$ between two ingestion rates is largely uniformly for $\mathrm{{}^{92}Mo}$$\mathrm{{}^{196}Hg}$. The average spread $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}$ between the different ingestion rates for the MLT and 3D-inspired scenarios is $1.22$, $1.58$, $1.64$, $1.78$, and $1.84\,\mathrm{dex}$. This shows that the non-linear and non-monotonic behaviour in Section III.2 is because of the changing location of peak burning in the convective-reactive environment and not just more material being present.
@@
Without ingestion of $\mathrm{C}$-shell material, the MLT, $1\times$, and $3\times$ scenarios see little to no production, but for the $10\times$ and $50\times$ scenarios there is a significant underproduction of the heavier $p$ nuclei. This is because in the fastest cases the species are advected to the bottom of the shell where the temperatures are purely destructive and there is no replenishment from the $\mathrm{C}$ shell. This shows that entrainment of $\mathrm{C}$-shell material is necessary for significant contributions from pre-explosive $\gamma$ process.
Another important feature of ingestion is that $n$-heavier isotopes are more favoured as the rate increases for all mixing speeds in our study. For isotopic pairs, production of the heavier isotope with respect to the lighter increases with ingestion rate due to the increase in released neutrons from burning $\mathrm{C}$ and $\mathrm{O}$. Figure 11 shows, for example, how $\mathrm{{}^{106,108}Cd}$ are similar for the no merger case but as ingestion rate increases $\mathrm{{}^{108}Cd}$ becomes increasingly more abundant compared to $\mathrm{{}^{106}Cd}$. Other isotopic pairs exhibit an increase of the heavier isotope when ingestion is present but $\mathrm{OP}$ between the two does not change with ingestion rate, such as $\mathrm{{}^{156,158}Dy}$.
@@
The sign of the ratio of isotopic pairs is largely preserved across the ingestion rates for the MLT, $1\times$, $3\times$, and $50\times$ scenarios, but the magnitude of the ratio does change. The few exceptions are $\mathrm{{}^{162,164}Er}$ in the $1\times$ and $3\times$ scenarios who become roughly equal for the fastest ingestion rate, $\mathrm{{}^{115}Sn}$ which becomes more abundant than $\mathrm{{}^{112,114}Sn}$ in the $3\times$ scenario. In the $10\times$ scenario production of isotopic pairs is largely equal for the slower ingestion rates, but at the fastest ingestion rate the ratio can become unequal. This shows that the the ingestion rate matters for the magnitude of the ratio between isotopic pairs, but that the sign of the ratio depends on which diffusion profile is used.
III.4 Impact from dips from convective quenching
@@
Here we present the impact of dips from convective quenching using the profiles shown in Figure 5 from Section II.2 which represent GOSH-like feedback and a partial merger. The results are shown in Figure 12.
@@
Figure 12: The overproduction compared to initial of the $p$ nuclei for the MLT, GOSH-like, and partial merger scenarios with ingestion rate $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.51\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount. Average $\mathrm{OP}$ for each scenario are provided as dashed lines and corresponds to the values presented in Table 1.
@@
The dip functions as a barrier for the convective-reactive flow, and can section off parts of the $\mathrm{O}$ shell. The partial merger scenario both limits the ingested $\mathrm{C}$-shell material and slows down the material from deeper in the $\mathrm{O}$ shell from advecting up to the top of the shell where very few reactions occur. The GOSH-like dip slows down the material from reaching their preferential temperatures, and also prevents material at the deepest part of the $\mathrm{O}$ shell from mixing up to the top of the shell which keeps it at hotter temperatures where it can be destroyed. This is also why the deeper dips suppress production more, as the velocities are slower by an additional factor of $10$ at the deepest point of the dip.
The suppression of production is largely uniform for all isotopes except for $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, $\mathrm{{}^{84}Sr}$, and $\mathrm{{}^{180}Ta}$. Additionally, it appears that isotopes $A\geq 138$ are more strongly affected by the stronger GOSH-like dip. The isotopes $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$ have a minor increase from these dips because the location of their production is at the bottom of the $\mathrm{O}$ shell where the dips are not present. The boost in production of $\mathrm{{}^{180}Ta}$ is because the peak production and destruction locations happen to be centered at the exact same location as the GOSH dip, which lowers its destruction and slightly boosts its production. Isotopic ratios are not significantly affected by the dips, although the magnitude of the ratios can increase slightly. This demonstrates the importance of both the location and magnitude of the convective dips in the $\mathrm{O}$ shell for convective-reactive $\gamma$ process.
III.5 Nuclear physics impact and mixing dependencies
@@
The MLT mixing scenario has an average spread in production of $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.56\;\mathrm{\,\mathrm{dex}}$, and the 3D-inspired scenarios have an average spread of $0.59\;\mathrm{\,\mathrm{dex}}$, $0.69\;\mathrm{\,\mathrm{dex}}$, $0.76\;\mathrm{\,\mathrm{dex}}$, and $0.79\;\mathrm{\,\mathrm{dex}}$ for the $1\times$, $3\times$, $10\times$, and $50\times$ scenarios respectively. We find that the spread in production increases with mixing speed because the material is able to reach hotter temperatures and the number of possible nucleosynthetic pathways changes.
@@
The spread of an individual isotope is dependent on the mixing scenario. Species like $\mathrm{{}^{106}Cd}$, $\mathrm{{}^{156}Dy}$, and $\mathrm{{}^{180}W}$ have a different spread in production for each mixing scenario. While the change can be monotonic for species like $\mathrm{{}^{180}W}$, $\mathrm{{}^{184}Os}$, and $\mathrm{{}^{190}Pt}$, for $\mathrm{{}^{106}Cd}$ and $\mathrm{{}^{112}Sn}$, and $\mathrm{{}^{130}Ba}$ they decrease for the $50\times$3D-inspired scenario.
@@
It is clearly important to consider the mixing conditions if an experiment is to be proposed to measure a reaction rate. As Table 2 shows for the $50\times$3D-inspired scenario, there are many reactions even for a single isotope that can be correlated uniquely in that mixing scenario. This clearly shows that a decreasing radial velocity profile and the exact magnitude of the mixing speeds are crucial for understanding the nuclear reactions in this convective-reactive environment.
There are also correlated reactions that all mixing scenarios share as shown in Table 3. Additionally, all downturn cases share correlations not found in the MLT scenario: $X(\mathrm{{}^{115}Sn})$ with $\mathrm{{}^{110}Sn}(\gamma,\alpha)$ and $X(\mathrm{{}^{138}Ce})$ with $\mathrm{{}^{138}Nd}(\gamma,p)$. However, the shared correlations are not of equal strength across all mixing scenarios. As an example the final mass fraction of $\mathrm{{}^{156}Dy}$ is correlated with $\mathrm{{}^{160}Er}$$(\gamma,\alpha)$, but in the $50\times$3D-inspired the correlation is much weaker. This underscores the possible difficulties in using 1D astrophysical sites to identify important reactions for nuclear physics experiments.
@@
$^{144}\mathrm{Sm}$ $^{196}\mathrm{Pb}(\gamma,n)$  $^{142}\mathrm{Sm}(\gamma,n)$  $^{142}\mathrm{Sm}(\gamma,p)$
$^{143}\mathrm{Sm}(\gamma,n)$
@@
[…]
Rauscher et al. (2016) studied the impact of nuclear uncertainties for the explosive $\gamma$-process for a $15\;\mathrm{\mathrm{M}_{\odot}}$ and two $25\;\mathrm{\mathrm{M}_{\odot}}$ models with solar metallicity using a Monte Carlo method. The spread in their 90% probability interval for the $15\;\mathrm{\mathrm{M}_{\odot}}$ model was $0.61\;\mathrm{\,\mathrm{dex}}$, $0.63\;\mathrm{\,\mathrm{dex}}$ for the $25\;\mathrm{\mathrm{M}_{\odot}}$ KEPLER model, and $0.99\;\mathrm{\,\mathrm{dex}}$ for the model from Hashimoto et al. (1989). This is comparable to the maximum spread found in this work, although we are considering pre-explosive $\gamma$ process. Rauscher et al. (2016) also provide tables of correlated rates, but only a handful of these rates appear in our Tables C1C5 including those they find with $r_{\mathrm{P}}\geq|0.65|$. This is because the convective-reactive environment we consider allows for different reaction pathways to be favoured depending on the mixing conditions.
IV Discussions and Conclusion
In this paper, we have shown that understanding the mixing details in the $\mathrm{O}$-burning shell during an $\mathrm{O}$-$\mathrm{C}$ shell merger is crucial for accurately modelling the nucleosynthesis of the $p$ nuclei. This work raises the importance of 3D hydrodynamic simulations for understanding the nucleosynthesis in $\mathrm{O}$-$\mathrm{C}$ shell mergers (Bazan & Arnett, 1994; Yadav et al., 2020; Rizzuti et al., 2024). We have demonstrated the convective-reactive nature of the nucleosynthesis in the $\mathrm{O}$ shell, where the timescales for advection and reaction are comparable and our results show that:
@@ IV Discussions and Conclusion
  • The ratio of isotopic pairs is sensitive to the mixing scenario.
@@
Although we find that the mixing conditions significantly impact the results of Section III.5, not all scenarios are equally likely to represent the conditions in a merger. The MLT mixing scenario, $1\times$, and $50\times$ case may not be representative of the conditions for realistic $\mathrm{O}$-$\mathrm{C}$ shell mergers. 3D hydrodynamic simulations show the $\mathrm{O}$ shell has a downturn to radial convective velocities and mixing speeds roughly 3103-10 times larger than what MLT predicts (Jones et al., 2017; Andrassy et al., 2020; Rizzuti et al., 2024). This suggests that the $3\times$ and $10\times$ downturn mixing scenarios are likely more representative of the conditions in a merger. The exploration done in this paper shows the importance of understanding the mixing conditions, but not all scenarios are equally likely to represent the conditions in a merger.
The results in this work are important for the interpretation of presolar grains. As shown in this work, the ratio of isotopic pairs is sensitive to the mixing conditions in the $\mathrm{O}$ shell. This means that comparing the grains data to the results of this work can be used as a diagnostic tool to constrain the mixing details of $\mathrm{O}$-$\mathrm{C}$ shell mergers and connect measured isotopic ratios to 3D hydrodynamics.
There are limitations to the results provided in this work and further extensions that could be done. We have focused on the impact of mixing conditions and varying reaction rates, but do not investigate how different stellar conditions are relevant to the nucleosynthesis or if changing the stable seeds from the $\mathrm{C}$ shell impact these results. The distribution of stable seeds and the metallicity of the star are of significant importance for the nucleosynthesis of the $p$ nuclei (Travaglio et al., 2015; Battino et al., 2020). It is possible that earlier in stellar evolution that weak $s$ process in the $\mathrm{C}$ shell could modify the stable seeds (Pignatari et al., 2010), which could be relevant for the lighter $p$ nuclei.
@@
Finally, $\mathrm{O}$-$\mathrm{C}$ shell mergers are crucial to the nucleosynthesis of a massive star prior to the CCSN regardless of explosive energy (Roberti et al., 2024). Even if the results in Figure 14 do not represent the whole of the nucleosynthesis of the $p$ nuclei, they are still crucial for understanding the nucleosynthesis in the $\mathrm{O}$ shell prior to the CCSN.
@@
The results of this work have implications beyond the $p$ nuclei. The light odd-Z elements $\mathrm{P}$, $\mathrm{Cl}$, $\mathrm{K}$, and $\mathrm{Sc}$ are also produced during $\mathrm{O}$-$\mathrm{C}$ shell mergers and based on our preliminary results are likewise impacted by varying the mixing conditions (Ritter et al., 2018a; Roberti et al., 2025). The long-lived radioactive isotope $\mathrm{{}^{40}K}$ ($t_{1/2}=1.25\times 10^{9}\;\mathrm{\mathrm{yr}}$), which is relevant to the heating of planets early in their formation (Frank et al., 2014; O’Neill et al., 2020), is also affected along with the stable $\mathrm{K}$ isotopes $\mathrm{{}^{39}K}$ and $\mathrm{{}^{41}K}$. Finally, observations have found $\mathrm{P}$-enhanced stars (Masseron et al., 2020; Brauner et al., 2023, 2024) which could be explained by a $\mathrm{O}$-$\mathrm{C}$ shell merger from a previous massive star.
@@
$\mathrm{O}$-$\mathrm{C}$ shell mergers have a potentially huge impact on galactic chemical evolution models (Ritter et al., 2018a), and massive star models show this feature regardless of metallicity and stellar evolution model (Roberti et al., 2025). Further work is needed to understand the impact of the macrophysical uncertainties in 1D stellar models on the nucleosynthesis of these light odd-Z elements.
@@
The authors would like to thank Stephen Mojzsis for the useful discussions. MP acknowledges the support to NuGrid from the “Lendulet-2023” Program of the Hungarian Academy of Sciences (LP2023-10, Hungary), the ERC Synergy Grant Programme (Geoastronomy, grant agreement number 101166936, Germany), the ERC Consolidator Grant funding scheme (Project RADIOSTAR, G.A. n. 724560, Hungary), the ChETEC COST Action (CA16117), supported by the European Cooperation in Science and Technology, and the IReNA network supported by NSF AccelNet (Grant No. OISE-1927130). MP also thanks the support from NKFI via K-project 138031 (Hungary). FH is supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant and acknowledges support from the NSERC award SAPPJ-797 2021-00032 Nuclear physics of the dynamic origin of the elements. We acknowledge support from the ChETEC-INFRA project funded by the European Union’s Horizon 2020 Research and Innovation programme (Grant Agreement No 101008324). This research has used the Astrohub online virtual research environment (https://astrohub.uvic.ca), developed and operated by the Computational Stellar Astrophysics group at the University of Victoria and hosted on the Digital Research Alliance of Canada Arbutus Cloud at the University of Victoria. This work benefited from interactions and workshops co-organized by The Center for Nuclear astrophysics Across Messengers (CeNAM) which is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-SC0023128.